5. Find the labor force participation rate and unemployment rate given the following information about an economy:

                Population (working age)   120

                Labor Force                          80        PLEASE JUST ANSWER THE QUESTION 6.

                # Employed                          70

  1. Assume in the problem above 5 of the unemployed became discouraged workers. Calculate new official unemployment rate in this economy. Comment on your finding.
Answer #1

Labor force = 80

Employed = 70

Labor force = Employed + Unemployed

80 = 70 + unemployed

Unemployed = 80 - 70

Unemployed = 10.

Initial unemployment rate

Unemployment rate = (Unemployed people / Labor force) * 100

Unemployment rate = (10 / 80) * 100

Unemployment rate = 12.5%

If 5 of the unemployed became discouraged worker, then the remaining number of unemployed would be 5. Discouraged worker is not counted in the Labor Force.

So, the labor force value will decrease from 80 to 75.

New unemployment rate:

Unemployment rate = (Unemployed people / Labor force) * 100

Unemployment rate = (5 / 75) * 100

Unemployment rate = 6.67%